OpenStudy (anonymous):

How do you answer 4-3x=x-8 in steps please.

OpenStudy (anonymous):

show us how far you can get,others have already replied to similar questions from you

OpenStudy (anonymous):

$$ 4-3x=x-8 \Rightarrow -4x = -12 \Rightarrow x =3 $$

OpenStudy (anonymous):

soo it would be. 3x+x=4+8 >>>>>>> 4x=12 >>>>>>> 12/4x = 3x

OpenStudy (anonymous):

4x=12 divide both sides by 4 4x/4=12/4 x=3